# Online Decision-Focused Learning

### v1

@@ 2 Framework
which is a regularized approximation of $w^{\star}_{t}$. When this surrogate is continuously differentiable—which holds for our choices of $\mathcal{R}$ in Section 4.1 and Section 4.2—we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine in place of $\nabla w^{\star}_{t}(\theta)$. This approach amounts to minimizing the surrogate
@@
instead of $f_{t}$, which admits gradients of the form
$\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
@@
Note that $\tilde{f}_{t}$ is not necessarily convex (even when $\theta\mapsto g(\theta,X)$ is linear) because it involves the parametrized minimizer $\tilde{w}_{t}(\theta)$. This prevents us from directly using known online convex optimization algorithms hazan2023introductiononlineconvexoptimization . However, a recent line of research kalai2005efficient ; agarwal2019learning ; suggala2020online ; xu2024online has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous—which is the case with our choices of regularizers in Section 4.1 and Section 4.2 (see Lemmas 2 and 7). These studies combine near-optimal oracles with perturbation techniques to establish sublinear bounds on the expected regret. In line with this literature, we also assume to have access to a near-optimal oracle. In the following assumption, we define for any $t\in[T]$, $\tilde{F}_{t}:\theta\mapsto\mathbb{E}[\tilde{f}_{t}(\theta)\,|\,\mathcal{H}_{t-1}]$.
H 3.
There exists $\xi>0$ such that for any $t\in[T]$, upon observing $\mathcal{H}_{t}$, we can compute $\vartheta_{t}\in\Theta$ satisfying almost surely:
$\tilde{F}_{t}(\vartheta_{t})\leqslant\inf_{\theta\in\Theta}\tilde{F}_{t}(\theta)+\xi\;.$ (7)
One way to design such oracles is to apply Stochastic Gradient Descent amari1993backpropagation to the differentiable function $\tilde{f}_{t}$. In this case, $\xi\geqslant 0$ captures both the error made by the offline minimization method generating $\vartheta_{t}$, and the gap between the local optimum to which it converges and the global optimum. In particular, $\xi$ vanishes for benign loss landscapes. Finally, note that in the case where $\tilde{w}_{t}(\theta)$ can only be approximated, then $\xi$ can also encompass the resulting error.
Leveraging the regularization and perturbation approaches previously mentioned, we now present our method, referred to as DF-OGD, in Algorithm 1.
@@
5:     Observe $\vartheta_{t}\in\Theta$ satisfying (7).
6:     Draw $\delta_{t}\sim\text{Unif}([0,1])$, compute $u_{t}=\vartheta_{t}+\delta_{t}(\theta_{t}-\vartheta_{t})$ and ~$\tilde{\nabla}_{t}(u_{t})=\nabla\tilde{w}_{t}(u_{t})^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
@@
At each time $t\in[T]$, Algorithm 1 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$. Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$ and computes the oracle $\vartheta_{t}$ from H 3. Next, it determines the gradient $\nabla\tilde{w}_{t}(u_{t})$, evaluated in a point $u_{t}$ drawn uniformly at random in $[\theta_{t},\vartheta_{t}]$. From there, it updates $\theta_{t+1}$ through the classical gradient step scheme of OGD, with the noticeable difference that the gradient is evaluated in $u_{t}$ instead of $\theta_{t}$.
Algorithm 1 requires determining $w^{\star}$, which can be efficiently achieved by classic numerical solvers in reasonable settings. However, we show in Appendix E that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}\right\rangle-\mathopen{}\mathclose{{}\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
Note that while DF-OGD relies on OGD, it is possible to use other algorithmic approaches such as Follow the Perturbed Leader (FTPL) suggala2020online , see DF-FTPL in Appendix B for which we provide a static regret analysis.
4 A theoretical study of convergence
We now provide convergence guarantees for Algorithm 1 in the form of expected dynamic regret bounds in two settings. First, we deal with the case where $\mathcal{W}$ is the simplex. Second, we consider the general setting where $\mathcal{W}$ is any bounded convex polytope with non-empty interior.
4.1 Online decision-focused learning on the simplex
We first focus on the case where $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, that is
[…]

### v2

@@ 2 Framework
which is a regularized approximation of $w^{\star}_{t}$. When this surrogate is continuously differentiable—which holds for our choices of $\mathcal{R}$ in Section˜3—we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine. This approach amounts to minimizing the surrogate
@@
instead of $f_{t}$, which admits gradients of the form $\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})$. Note that there is a natural trade-off in the choice of the regularization parameter $\alpha_{t}$. On the one hand, choosing a large $\alpha_{t}$ makes the function $\tilde{w}_{t}(\theta)$ smoother. On the other hand, the larger $\alpha_{t}$ is, the more $\tilde{w}_{t}$ deviates from the true function $w^{\star}_{t}$ that we aim to approximate. As we will see later, it is possible to balance these two extremes by carefully tuning the parameter $\alpha_{t}$.
On the choice of regularization on a general polytope.
In what follows, we write $\mathcal{W}$ as the intersection of $n$ half-spaces, that is $\mathcal{W}=\{w\in\mathbb{R}^{d}:\>A^{\mathsf{\scriptscriptstyle T}}w-b\preccurlyeq 0\}$ where $n>0$ is the number of faces, $A\in\mathbb{R}^{d\times n}$ and $b\in\mathbb{R}^{n}$. We assume that $\mathcal{W}$ is not degenerated, i.e. $AA^{\mathsf{\scriptscriptstyle T}}\in\mathbb{R}^{d\times d}$ is full-rank.
It remains to determine what regularizer $\mathcal{R}$ to choose in (6). We recall that our aim is to obtain a $\tilde{w}_{t}(\theta)$ in (5) which is differentiable for any $\theta\in\Theta$. A possible strategy is to choose $\mathcal{R}$ so $\tilde{w}_{t}$ remains in the strict interior of $\mathcal{W}$. In this case, $\tilde{w}_{t}(\theta)$ is differentiable in a neighborhood of any $\theta\in\Theta$ and admits a close-form Jacobian $\nabla\tilde{w}_{t}(\theta)$ by the implicit function theorem. A natural choice to force $\tilde{w}_{t}$ to remains in the interior of $\mathcal{W}$ is the corresponding log-barrier function:
$\mathcal{R}:w\mapsto-\sum_{i=1}^{n}\ln(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}w)\;,$ (7)
where $A_{i}\in\mathbb{R}^{d}$ is the $i$-th column of $A$. With this choice of regularization, we show in Lemma˜4 that $\nabla\tilde{w}_{t}$ has an explicit formulation, allowing its use in practice.
Remark 1.
We remark that in the special case where $\mathcal{W}=\{w\in\mathbb{R}^{d}:w\succcurlyeq 0\;\text{,}\;{\bf 1}^{\mathsf{\scriptscriptstyle T}}w=1\}$ is the simplex of $\mathbb{R}^{d}$, an alternative choice for $\mathcal{R}$ in (5) is the negative entropy $\mathcal{R}_{0}:w\mapsto\sum_{i\in[d]}w_{i}\ln(w_{i})$. In this case, $\tilde{w}_{t}(\theta)$ in (5) reduces to the softmax mapping, which is differentiable and admits a close-form Jacobian. In Appendix˜C, we theoretically study the performances of our algorithms in this special case.
Approximate oracles to handle non-convexity.
@@
The bi-level structure of the problem yields unexpected properties. In particular, even when the regularizer $\mathcal{R}$ is strongly-convex in (26)—which the case with both log-barriers and negative entropy—and the model $g$ is simple (such as $g(\theta,X)=X\theta$), $\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{\left\langle\tilde{w}_{t}(\theta),\bar{g}_{t}(X_{t})}}\right\rangle$ may not be convex, but is lipschitz (see Lemmas˜5 and 1). This prevents us from directly using known online convex optimization algorithms [37]. However, a recent line of research [3, 67, 75] has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous—which is the case with log-barriers or negative entropy (see Lemmas˜1 and 5). These studies combine near-optimal oracles with perturbation techniques to establish sublinear bounds on the expected regret. In line with this literature, we also assume to have access to an approximate offline optimization oracle, a notion that appeared in a slightly modified form in Suggala and Netrapalli [67].
Definition 1.
An $\xi$-approximate offline optimization oracle (or approximate oracle), adapted to a class $\mathcal{C}$ is a mapping $\mathbf{O}_{\xi}$ taking a function $f\in\mathcal{C}$ and outputting $\vartheta=\mathbf{O}_{\xi}(f)\in\Theta$ satisfying:
$f(\vartheta)\leqslant\inf_{\theta\in\Theta}f(\theta)+\xi\;.$ (8)
The notion of an approximate oracle formalizes that, in non-convex settings, we cannot rely on subroutines that provably attain a global minimizer of $\tilde{f}_{t}$ at time $t$ (as offline gradient descent would in convex problems). Instead, we must settle for local minimizers, whose quality is characterized by a parameter $\xi$—which vanishes in favorable loss landscapes. As a concrete example, consider the stochastic gradient descent (SGD) algorithm as $\mathbf{O}_{\xi}$. A large body of work has shown that, even when $\Theta$ is a class of deep neural networks, SGD can converge to local minimizers [31, 55, 58, 20], justifying that $\mathbf{O}_{\xi}$ is indeed an approximate oracle. A more detailed discussion is provided in Appendix˜B.
We conclude this remark by mentioning that it is possible to conduct a non-convex analysis without an approximate oracle by considering the weaker notion of local regret defined as $\sum_{t\in[T]}\mathopen{}\mathclose{{\left\|\nabla F_{t}(\theta_{t})}}\right\|$; see [38, 6, 80, 36]. However, such a definition is not meaningful in the DFL framework, where gradients are zero or undefined due to the structure of the problem.
The combined use of approximate oracles and regularization allows us to derive original online algorithms. In Section˜3.1, we focus on a variant of the FTL algorithm, enjoying a static regret bound, while in Section˜3.2 we propose a version of OGD enjoying a dynamic regret bound, both results holding in the non-convex case.
3.1 Decision-Focused Follow The Perturbed Leader
Our first algorithm is displayed in Algorithm˜1, where $\mathrm{Exp}(\eta)$ refers to an exponential distribution with parameter $\eta>0$. It is inspired from the classic Follow-the-Leader approach [45], which consists in making a decision minimizing the sum of objective functions observed so far. When losses are non-convex, it is common to inject random noise at each step to regularize the total cost function, resulting in an approach known as Follow-the-Perturbed-Leader [43, 67]. We employ this strategy as $\mathsf{Alg}_{t}$ to update $\theta_{t+1}$ in Algorithm˜1. Note that the oracle in line 6 of Algorithm 1 crucially works with the regularized losses $\tilde{f}_{1},\ldots,\tilde{f}_{t}$. This allows to use in practice gradient-based methods to obtain an approximate minimizer as per our discussion in Section˜3, since these surrogate losses are differentiable and Lipschitz.
Algorithm 1 Decision-Focused Follow The Perturbed Leader DF-FTPL
1:Input: horizon $T>0$, initialization $\theta_{1}\in\Theta$, $\xi$-approximate oracle $\mathbf{O}_{\xi}$ and history $\mathcal{H}_{0}=\emptyset$.
2:for each $t\in\{1,\ldots,T\}$ do
5:  Draw $\sigma_{t}\in\mathbb{R}^{d}$ such that for all $j\in[d]$, the $j$-th component $\sigma_{j,t}\sim\mathrm{Exp}(\eta)$.
6:  Update
$\theta_{t+1}=\mathbf{O}_{\xi}\mathopen{}\mathclose{{\left(\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{\left\langle\sigma_{t},\cdot}}\right\rangle}}\right)$
7:end for
Note that Algorithm˜1 requires determining $w^{\star}$, which can be efficiently achieved by classic numerical solvers in reasonable settings. However, we show in Appendix˜F that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
We provide a theoretical guarantee on the convergence of Algorithm˜1 in the form of the following static regret bound.
Theorem 1.
Assume H˜1, H˜2 and having access to an $\xi$-approximate oracle $\mathbf{O}_{\xi}$ adapted to $\{\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{\left\langle\sigma_{t},\cdot}}\right\rangle\}_{t\in[T]}$. Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-FTPL (Algorithm˜1) instantiated with learning step $\eta>0$ and regularization coefficients $\alpha_{t}=\alpha>0$ for any $t$. Then:
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha n}}\right)\;,$
Furthermore, taking $\eta\propto m^{1/4}T^{-3/4}n^{-1/2}$ and $\alpha\propto m^{3/4}n^{1/2}T^{-1/4}$ yields:
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(m^{3/4}\sqrt{n}T^{-1/4}+\xi}}\right).$
The proof of Theorem˜1 can be found in Section˜G.2. In particular, Theorem˜1 shows that DF-FTPL enjoys an average regret bound decaying in $T^{-1/4}$ as long as $\xi=\mathcal{O}(T^{-1/4})$. While it features a polynomial dependency $m^{3/4}$ on the dimension of $\Theta$, it only depends on the dimension of the decision space $\mathcal{W}$ through a $\ln\ln(d)$ term. This makes Algorithm˜1 a particularly competitive approach when the decision space is of high dimension.
It is informative to compare our guarantee to existing bounds in the literature. [67], who also study non-convex online learning, achieves a rate of $\mathcal{O}(T^{-1/2})$ as long as their offline oracle satisfies $\xi=\mathcal{O}(T^{-1/2})$, at the cost of a degraded $m^{3/2}$ dependency on the dimension. Their faster rate in $T$ comes from the fact that they tackle a simpler, single-level problem as compared to our bi-level, non differentiable setting. Indeed, we need to regularize the inner problem to overcome non-differentiability as discussed in Section˜3. This introduces an additional trade-off on the regularization strength $\alpha$ on top of the usual trade-off in the learning rate $\eta$, which is reflected in our rate.
3.2 Decision-Focused Online Gradient Descent
While DF-FTPL (Algorithm˜1) benefits from a converging static regret bound, the techniques involved in [67] are not enough to reach dynamic regret guarantees, which is particularly relevant in highly non-stationary environments where the optimal decisions may change significantly from one round to another. To this end, we go beyond the Follow-the-Leader approach and propose an original algorithm based on the celebrated Online Gradient Descent [81]. This procedure, which we call DF-OGD, is presented in Algorithm˜2.
In words, at each time $t\in[T]$, Algorithm˜2 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$. Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$ and computes one of its near minimizer via the offline oracle from Definition˜1. Next, to handle non-convexity, it determines the gradient $\nabla\tilde{w}_{t}(u_{t})$, evaluated in a point $u_{t}$ drawn uniformly at random in $\theta_{t},\vartheta_{t}]$. From there, it updates $\theta_{t+1}$ through the classical gradient step scheme of OGD.
[…]
Moreover, note that Algorithm˜1 is instanciated with a single pair of parameters $(\alpha,\eta)$ whereas Algorithm˜2 uses on a sequence $(\alpha_{t},\eta_{t})_{t\in[T]}$. This additional flexibility is crucial for the algorithm to adapt to the variation of the problem so as to maintain a low dynamic regret.
@@ 3.2 Decision-Focused Online Gradient Descent
5:  Get $\vartheta_{t}=\mathbf{O}_{\xi}(\tilde{f}_{t})$.
6:  Draw $\delta_{t}\sim\text{Unif}([0,1])$, compute $u_{t}=\vartheta_{t}+\delta_{t}(\theta_{t}-\vartheta_{t})$ and $\tilde{\nabla}_{t}(u_{t})=\nabla\tilde{w}_{t}(u_{t})^{\mathsf{\scriptscriptstyle T}}\bar{g}_{t}(X_{t})\;.$
@@
We provide a convergence guarantee for Algorithm˜2 with the following dynamic regret bound.
Theorem 2.
Assume H˜1, H˜2, access to a $\xi$-approximate oracle adapted to $\{\tilde{f}_{t}\}_{t\in[T]}$. Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-OGD (Algorithm˜2) instantiated with the non-increasing sequence $\mathopen{}\mathclose{{\left(\eta_{t}}}\right)_{t\in[T]}$ and regularization coefficients $(\alpha_{t})_{t\in[T]}$. Then:
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\mathbb{E}\mathopen{}\mathclose{{\left[\frac{1+P_{T}}{T\eta_{T}}+\frac{1}{T}\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+n\alpha_{t}}}\right]+\xi}}\right)\;.$
where $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|$, $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-OGD and $\tilde{\mathcal{O}}$ contains polynomial dependency in $\ln(1/\alpha),\ln(\ln(d))$.
Furthermore, assume $(t^{-1}(1+P_{t}))_{t\geqslant 1}$ is non-increasing almost surely with $P_{t}=\sum_{s=1}^{t}\mathopen{}\mathclose{{\left\|\vartheta_{s+1}-\vartheta_{s}}}\right\|$. Then, using $\alpha_{t}\propto n^{-1/2}t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto n^{-1/2}t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to :
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\mathbb{E}\mathopen{}\mathclose{{\left[\sqrt{n}(1+P_{T})^{1/4}T^{-1/4}}}\right]+\xi}}\right)\;.$
The complete statement of the main bound, including constants is provided in Section˜G.3. Our dynamic regret bound naturally depends on $P_{T}$, which captures the problem’s variability by measuring the cumulative distance between approximate minimizers over time. When $\xi=\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$, the average dynamic regret decreases at the rate $\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$. Notably, the bound is independent of the dimension of $\Theta$, and depends only mildly on the dimension of $\mathcal{W}$ through a $\ln\ln(d)$ factor. This constitutes a key strength of the optimistic strategy underlying DF-OGD, making it particularly well-suited for high-dimensional settings.
It is instructive to compare our guarantee with those established in recent studies on dynamic regret. For instance, [77] derive a $\mathcal{O}(\sqrt{T(P_{T}+1)})$ bound. However, their setting is not directly comparable to ours, as they consider a simpler single-level problem with differentiable and convex objectives. In contrast, our framework involves non-convex, non-differentiable losses due to the bi-level nature of decision-focused learning. More recently, [41] obtained a $\mathcal{O}((1+P^{\infty}_{T})^{1/3}T^{-1/3})$ bound, where $P^{\infty}_{T}:=\sum_{t=1}^{T}\|f_{t+1}-f_{t}\|_{\infty}$. Yet, this rate is achieved under substantially more favorable conditions: a single-level problem with losses that are strongly convex or Lipschitz, an additional assumption of “quasi-stationary” and a more challenging path involving the full landscapes of $f_{t}$s. By comparison, our losses are neither Lipschitz (indeed, they may even be discontinuous due to the linearity of the lower-level problem) nor required to be stationary over time—the only restriction being H˜2 to hold. We view the ability of our algorithm to achieve efficient convergence despite these demanding conditions as a core contribution of our work.
4 Experiments
In this section, we compare the performances of our algorithms DF-FTPL and DF-OGD to two important benchmarks, namely prediction-focused learning and SPO [25].
Setting.
Our experimental setup is inspired by the knapsack example from Mandi et al. [54]. More precisely, we consider a decision maker who must pick at each $t\in[T]$ an object $v_{t}\in\mathcal{V}$ among $K$ items denoted $\mathcal{V}=\{1,2,\ldots,K\}$ with respective costs $\bar{g}_{t}(X)=(\bar{g}_{t,1}(X),\ldots,\bar{g}_{t,K}(X))\in[0,1]^{K}$ depending on some covariates $X\in\mathcal{X}$. At the beginning of each period, the decision-maker only observe covariates $X_{t}\in\mathcal{X}\subset\mathbb{R}^{p}$, and have at their disposal a parametric model $g:\Theta\times\mathcal{X}\rightarrow[0,1]^{K}$ to predict $\bar{g}_{t}$. Given their current parameter $\theta_{t}\in\Theta$, they predict item costs as $g(\theta_{t},X_{t})=(g_{1}(\theta_{t},X_{t}),\ldots,g_{K}(\theta_{t},X_{t}))$ and pick an item $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}g_{i}(\theta_{t},X_{t})$. This setting is depicted in Figure˜2 for $K=2$ in Appendix˜E. After having made their decision, the decision-maker observes the true item costs $\bar{g}_{t}(X_{t})$, and update $\theta_{t+1}$ for the next round based on this feedback. Note that this setting can directly be mapped in the simplex example of Remark˜1, since $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}_{0}}\mathopen{}\mathclose{{\left\langle w,g(\theta_{t},X_{t})}}\right\rangle$, $\mathcal{W}_{0}$ being the simplex of $\mathbb{R}^{K}$. For this reason, we instantiate our algorithms DF-OGD and DF-FTPL with the negative entropy regularizer, following Remark˜1.
Synthetic data.
We instantiate the previous problem with the following synthetic data. For any $t\in[T]$, we draw $X_{t}\in\mathbb{R}^{K\times p}$ with correlated rows, and generate a cost vector $c_{t}(X_{t})\in\mathbb{R}^{k}$ as :
$c_{t}(X_{t})=A\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon_{t}\;,$ (9)
where $A>0$, $\varepsilon_{t}\sim\mathrm{N}(0,I_{K})$ is a Gaussian noise and $\theta^{\star}_{t}\in\mathbb{R}^{p}$ is a parameter which satisfies $\theta^{\star}_{t}=\nicefrac{{1}}{{2}}\,\theta^{\star}+\nicefrac{{1}}{{2}}\,\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{p})$ for some $\theta^{\star}\in\mathbb{R}^{p}$. This is a challenging data generating process, since $\bar{g}_{t}(X_{t})=\sin^{4}((2X_{t}^{\mathsf{\scriptscriptstyle T}}\theta^{\star}_{t})^{-1})$ is non-stationary and highly non-linear, and features are correlated. Equation˜9 is discussed more in detail in Appendix˜E. To predict $c_{t}$ from $X_{t}$, we assume that the decision-maker has access to a class of linear predictors of the form $g:(\theta,X)\mapsto X\theta$.
Benchmarks.
In this setting, we compare the performances of DF-FTPL (Algorithm˜1) and DF-OGD (Algorithm˜2) to two benchmarks. First, we implement Prediction-Focused Online Gradient Descent (PF-OGD). This strategy consists in training in an online manner the model $g$ at each timestep so it minimizes the statistical loss $\ell^{\mathrm{mse}}:(v,X\theta)\mapsto\mathopen{}\mathclose{{\left\|v-X\theta}}\right\|^{2}$, irrespective of the downstream decision problem. Then, decisions are greedily made based on the predictions of the model. This approach is formally described in Algorithm˜3 in Appendix˜E. Second, we compare our algorithms to an online version of the Smart Predict-then-Optimize (online SPO, see Algorithm˜4) approach from [25]. This algorithm introduces a differentiable and convex loss which upper-bounds the actual decision-focused loss. Given its effectiveness and versatility, it is considered as a very important benchmark in the literature.
Results.
In Figure˜1, we plot on the left-hand panel the average cumulated cost $t\mapsto t^{-1}\sum_{s=1}^{t}\bar{g}_{s,v_{s}}(X_{s})$ incurred by DF-OGD, DF-FTPL, PF-OGD and online SPO over 10 runs of $T=5000$ timesteps, as well as the associated 95% confidence intervals. On the right-hand panel, we plot the average Mean Square Error (MSE) resulting from the sequence of prediction parameters $(\theta_{t})_{t\in[T]}$. It appears that both DF-FTPL and DF-OGD outperform PF-OGD and online SPO from a decision point of view, while incurring a higher MSE. This is in line with decision-focused approach, which cares about decision cost rather than statistical loss. This experiment shows that (i) DFL outperforms PFL when models are clearly mispecified and (ii) our algorithms also outperforms the celebrated online SPO.
Figure 1: Average cumulated cost (left) and prediction errors (right) of DF-OGD, PF-OGD, DF-FTPL and online SPO.
We also mention the presence of additional numerical experiments in Appendix˜E, to study how model misspecification affects the relative performances of DFL and PFL.
5 Conclusion.
Decision-focused learning offers a promising way to integrate prediction into decision making. We extend its analysis from the batch to the online setting, enabling non-stationary data and varying objectives. Our algorithms DF-FTPL and DF-OGD comes with provable guarantees, and empirically outperform prediction-focused learning and online SPO in our experiment.
We believe that this work can be extended in several ways. First, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operators. Second, it would be valuable to investigate less adversarial environments, such as those involving i.i.d. data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge. Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical perspective.
Appendix A H˜2 is satisfied in the Gaussian linear case.
In this section, we prove that H˜2 holds true when $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, $g$ is a linear model and $X$ has columns distributed according to a Gaussian distribution. In what follows, we denote by $X_{j}\in\mathbb{R}^{d}$ the $j$-th column of $X$.
Proposition 1.
Assume that $g:(\theta,X)\in\Theta\times\mathcal{X}\mapsto X\theta$ where $\Theta=\{\theta\in\mathbb{R}^{m}:\>\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}=1\}$ and $X=(X_{1}\,|\,\ldots\,|\,X_{m})\in\mathbb{R}^{d\times m}$ where $X_{j}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(\bm{0},\bm{I})$. Then, H˜2 is true for any $\theta\in\Theta$ with
$C_{0}=\frac{d(d-1)}{2\sqrt{\pi}}\quad\text{and}\quad\beta=1\;.$
Proof.
Let $\theta\in\Theta$. In what follows, we denote $i=I(\theta)$ where $I(\theta)$ is defined as in H˜2. Since $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, $v_{j}$ is the $j$-th vector of the canonical basis of $\mathbb{R}^{d}$. For any $\varepsilon>0$ we have:
$\mathbb{P}\mathopen{}\mathclose{{\left(\inf_{j\neq i}{\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle X\theta,v_{j}}}\right\rangle-\mathopen{}\mathclose{{\left\langle X\theta,v_{i}}}\right\rangle}}\right|}\,\leqslant\,\varepsilon}}\right)$ $\leqslant\mathbb{P}\mathopen{}\mathclose{{\left(\bigcup_{j\neq i}\,\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle X\theta,v_{j}}}\right\rangle-\mathopen{}\mathclose{{\left\langle X\theta,v_{i}}}\right\rangle}}\right|\,\leqslant\,\varepsilon}}\right\}}}\right)$
$\leqslant\sum_{j\neq i}\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\theta,X^{\mathsf{\scriptscriptstyle T}}(v_{j}-v_{i})}}\right\rangle}}\right|\,\leqslant\,\varepsilon}}\right)$
$=\sum_{j\neq i}\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\theta,X_{j}-X_{i}}}\right\rangle}}\right|\leqslant\varepsilon}}\right)\;.$ (10)
Since for any $j\neq i$, $\mathopen{}\mathclose{{\left\langle\theta,X_{j}-X_{i}}}\right\rangle\sim\mathcal{N}(0,2\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}^{2})$ with density $f_{j}$, we have:
$\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\theta,X_{j}-X_{i}}}\right\rangle}}\right|\,\leqslant\,\varepsilon}}\right)$ $=\int_{-\varepsilon}^{\varepsilon}f_{j}(x)\,\mathrm{d}x=\mathopen{}\mathclose{{\left(2\sqrt{\pi}\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}}}\right)^{-1}\int_{-\varepsilon}^{\varepsilon}\exp\mathopen{}\mathclose{{\left(-\frac{x^{2}}{4\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}^{2}}}}\right)\,\mathrm{d}x$
and since the integrand reaches its maximum in $x=0$,
$\leqslant\varepsilon\mathopen{}\mathclose{{\left(\sqrt{\pi}\mathopen{}\mathclose{{\left\|\theta}}\right\|_{2}}}\right)^{-1}\;.$ (11)
Then, plugging (11) in (10) and summing over all pairs $(i,j)$ such that $j\neq i$ yields the desired result. ∎
Note that the $d^{2}$ factor in $C_{0}$ comes from the use of an union bound, and could be reduced with a refined analysis.
Appendix B Extended discussion on the relevance of approximate oracles.
The notion of an approximate oracle reflects the fact that, in non-convex settings, we cannot rely on subroutines that provably reach a global minimizer of $\tilde{f}_{t}$ at time $t$ (as offline gradient descent would in convex problems). Instead, we must settle for local minimizers, whose quality is governed by a parameter $\xi$—which vanishes in favorable loss landscapes.
As a concrete example, consider $\mathbf{O}$ as the stochastic gradient descent (SGD) algorithm. Even in the context of deep networks, it is plausible that $\mathbf{O}$ converges to a local minimizer. Indeed, a large body of work has analyzed SGD in non-convex settings, establishing convergence to stationary points either in expectation [31] or almost surely [55, 58, 20]. Such stationary points may correspond to local/global minima or saddle points.
Recent studies further show that SGD avoids saddle points. For instance, Mertikopoulos et al. [55] proved that the trajectories of SGD almost surely avoid all strict saddle manifolds—i.e., sets of critical points $x$ where the Hessian has at least one negative eigenvalue. These manifolds include connected families of non-isolated saddle points, a phenomenon common in the loss landscapes of overparametrized neural networks [48].
Beyond SGD, similar guarantees extend to more general methods. In particular, the stochastic Riemannian Robbins–Monro method (a broad template encompassing various algorithms) has been shown to converge almost surely to local or global minima [40].
Appendix C Supplementary results on the simplex.
We now provide convergence guarantees for Algorithms˜1 and 2 when $\mathcal{W}$ is the simplex. To do so we choose another regularizer than done in the main document.
C.1 Additional framework
We focus on the case where $\mathcal{W}$ is the simplex of $\mathbb{R}^{d}$, that is
$\mathcal{W}_{0}=\mathopen{}\mathclose{{\left\{w\in\mathbb{R}^{d}:w\succcurlyeq 0,\;{\bf 1}^{\mathsf{\scriptscriptstyle T}}w=1}}\right\}\;.$